OpenStudy (anonymous):

What is 11 2/3 in radical notation?

OpenStudy (anonymous):

radical notation is when there is a radical in the answer... no... that would not be in radical notation...

OpenStudy (anonymous):

was i right @SolomonZelman

OpenStudy (solomonzelman):

it looks to me as though, that you need this.\[\huge\color{blue}{ a^{-b}=\frac{1}{a^b} }\] Not a fraction, a fractional exponent into a radical form.

OpenStudy (anonymous):

I am wondering do you mean 11 to the 2/3 power?

OpenStudy (solomonzelman):

Oh, you need this, my bad, here \[\huge\color{blue}{ a^ \frac{b}{c}=\sqrt[c]{a^b} }\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{SO,~~~~~~~~ 11^\frac{2}{3} =?}\]